OpenStudy (anonymous):

Using the given points, determine the slope. (0, 32) and (100,212) Slope = -9/5 Slope = -5/9 Slope = 5/9 Slope = 9/5

OpenStudy (alexandervonhumboldt2):

use slope formula: \[\frac{ y_{1}-y_{2} }{ x_{1}-x_{2} }\]

OpenStudy (alexandervonhumboldt2):

\[\frac{ y_{1}-y_{2} }{ x_{1}-x_{2} }\]=\[\frac{ 32-212 }{ 0-100 }=solve\]

OpenStudy (anonymous):

Ok can you give me a min to solve?

OpenStudy (alexandervonhumboldt2):

yeah

OpenStudy (anonymous):

I got 1.8 but I can't figure out how to put it into a fraction

OpenStudy (alexandervonhumboldt2):

look: you have -180/-90=180/90=2

OpenStudy (anonymous):

slope formula is y2- y1/ x2-x1

OpenStudy (alexandervonhumboldt2):

@Pandalova96

OpenStudy (anonymous):